3. Test the claim that a population mean exceeds 40. You have a sample of 40 items for which the sample mean x -42 and sample standard deviation s 7. Use significance level α 0.05. With worK please
Answer #1

Answer

we have population = μ = 40

sample mean = ar{x} = 42

sample standard deviation = s = 7

sample size = n = 40

we have to test the claim that the population mean exceeds 40, i.e. mu is greater than 40. So, it is a right tailed hypothesis.

First, we will calculate the test statistics and then we will find its corresponding p value to conclude about the claim.

Null hypothesis:- μ = 40 (Assuming that the population mean is equal to 40)

Alternate hypothesis:- μ>40 (Assuming that the population mean is greater than 40)

We will use t distribution because population standard deviation is unknown.

So, formula for t test statistic is given as  

t = (ar{x}-mu)/(s/sqrt{n})

setting the values, we get

t(42 40)/(7/v40)

this gives us

t (42-40)/(7/V40) 2/1.11-1.81

Degree of freedom = n-1 = 40-1 = 39

now, we have to find the p value corresponding to t value of 1.81 and degree of freedom 39

using the t distribution table, we get

p value= 0.0390

So, it is clear that p value is less than 0.05 level of significance.

Thus, we can say that the result is significant and we can reject the null hypothesis.

We can conclude that there is enough evidence to support the claim that the mean value exceeds 40